OpenStudy (anonymous):

The expression 700(1.005644)4x shows the quarterly interest rate Letrice earns after x years. Which expression shows the x isolated in the exponent in an equivalent form? 700(0.251411)x 700(4.022576)x 700(1.022768)x 700(1.001408)x

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

yes

OpenStudy (anonymous):

can you help me with this question

OpenStudy (dan815):

ok u know how 4*3*2*1 = 1*2*3*4 or any other way you , you want to multiply those numbers together

OpenStudy (anonymous):

24

OpenStudy (dan815):

yeah im just talking about the idea here they gave you 4 terms aswell in diguise

OpenStudy (anonymous):

ok so 24

OpenStudy (dan815):

700* 1.005644 * 4x =700*(1.005644 * 4) *x

OpenStudy (dan815):

what will it looks like when you multiply the 4 to that number instead of x

OpenStudy (anonymous):

yea i dont know ... im 100% lost

OpenStudy (dan815):

what is 1.005644 * 4

OpenStudy (anonymous):

4.02576

OpenStudy (dan815):

700* 1.005644 * 4x =700*(1.005644 * 4) *x = 700*4.02567 *x

OpenStudy (dan815):

see how that works?

OpenStudy (anonymous):

yes